Let X and Y be independent random variables with X = N(0, 1) and Y = Exp(1). Find E( |X| (Y + 1)^2 ).
Answer #1

solution- Here we find the E((x) * (Y+1²) Given dates. .X = N(0.1) Y = EXP (1) tere X and Y are independent random variables.E (W) = a ( E[(9*v*]. = E [2(3+)] - E (04) + E (3) =2E(4) + E(2) = 20+E(2) F: E(y) = 7 2x1 of 2 46(2) 22 . E [(y+1)] = 4 aE [